# Deep Learning-based MRI Reconstruction with Artificial Fourier Transform Network (AFTNet)

### v2

@@ 6 Conclusion
In conclusion, we propose AFT, a novel artificial Fourier transform framework that determines the mapping between k-space and image domain as conventional DFT while having the ability to be fine-tuned/optimized with further training. The flexibility of AFT allows it to be easily incorporated into any existing deep learning network as learnable or static blocks. We then utilized AFT to design our AFTNet, which implements complex-valued U-Net to extract features in the k-space and/or image domain. We aim to combine reconstruction and acceleration/denoising tasks into a unified network that simultaneously enhances the image quality by removing artifacts directly from the k-space and/or image domain. The proposed methods are evaluated on datasets with additional artifacts, different contrasts, and different modalities. Our AFTNet achieves competitive results compared with other methods and proves to be more robust to noise and contrast differences. An extensive study on various system fields, various modalities, and various tasks demonstrates the effectiveness and generality of AFTNet.

### v3

@@ 7 Conclusion
In conclusion, we propose AFT, a novel artificial Fourier transform framework that determines the mapping between k-space and image domain as conventional DFT while having the ability to be fine-tuned/optimized with further training. The flexibility of AFT allows it to be easily incorporated into any existing deep learning network as learnable or static blocks. We then utilized AFT to design our AFTNet, which implements complex-valued UNet to extract features in the k-space and/or image domain. We aim to combine reconstruction and acceleration/denoising tasks into a unified network that simultaneously enhances the image quality by removing artifacts directly from the k-space and/or image domain. The proposed methods are evaluated on datasets with additional artifacts, different contrasts, and different modalities. Our AFTNet achieves competitive results compared with other methods and is found to be robust to noise differences. An extensive study on various system fields, various modalities, and various tasks demonstrates the effectiveness and generality of AFTNet. Our current claims are supported by quantitative and objective visual analyses. In future works, neuroradiologist evaluations could be incorporated to further strengthen the qualitative assessment.
Appendix A
System
Model
Field
Strength (T)
Institution
Name
Protocol
Name
Encoded
Matrix Size
Field of View
TR
TE
Sequence
Type
Count
Avanto 1.494 NYU AX 640 x 272 x 1 440 x 186.9725 x 7.5 5120 103 TurboSpinEcho 1
Avanto 1.494 NYU AX 640 x 320 x 1 440 x 220 x 7.5 4000 107 TurboSpinEcho 1
Avanto 1.494 NYU AX 640 x 320 x 1 440 x 220 x 7.5 5120 103 TurboSpinEcho 373
Avanto 1.494 NYU AX 640 x 320 x 1 440 x 220 x 7.5 5120 107 TurboSpinEcho 295
Avanto 1.494 NYU Medical Center AX T2_FBB 640 x 320 x 1 440 x 220 x 7.5 5120 103 TurboSpinEcho 1
TrioTim 2.8936 NYU AX 768 x 308 x 1 440 x 176.5392 x 7.5 6000 107 TurboSpinEcho 1
TrioTim 2.8936 NYU AX 768 x 350 x 1 440 x 200.4446 x 7.5 6000 107 TurboSpinEcho 1
TrioTim 2.8936 NYU AX 768 x 392 x 1 440 x 224.62 x 7.5 6000 107 TurboSpinEcho 141
TrioTim 2.8936 NYU AX 768 x 392 x 1 460 x 234.83 x 7.5 6000 107 TurboSpinEcho 14
TrioTim 2.8936 NYU Clinical Cancer Ctr AX T1 PRE_FBB 640 x 260 x 1 440 x 178.75 x 5 264 2.88 Flash 1
TrioTim 2.8936 NYU Clinical Cancer Ctr AX T1 PRE_FBB 640 x 320 x 1 440 x 220 x 5 250 2.88 Flash 1
TrioTim 2.8936 NYU Clinical Cancer Ctr AX T1 PRE_FBB 640 x 320 x 1 440 x 220 x 5 264 2.88 Flash 106
TrioTim 2.8936 NYU Clinical Cancer Ctr AX T1 PRE_FBB 640 x 320 x 1 460 x 230 x 5 264 2.88 Flash 5
TrioTim 2.8936 NYU Clinical Cancer Ctr AX T1 PRE_FBB 640 x 320 x 1 480 x 240 x 5 264 2.88 Flash 2
Table 1: Imaging parameters. The name AX indicates T2w imageing protocol.
Figure 1: 1D 4x equal-spaced sampling mask with 8% of low-frequency columns retained. While space indicates the signal retained and black space indicates the signal masked out.